Word Problem for equadratic equations and inequalities.
An architect has designed a motel pool within a rectangular area that is fenced on three sides. If she uses 60 yards of fencing to enclose an area of 352 square yards, then what are the dimensions for L and W. Assume L is greater than W

Answer by checkley71(8403)

L+2W=60
L=60-2W NOW REPLACE L BY(60-2W) IN THE FIRST EQUATION & SOLVE FOR W
352=(60-2W)W
352=60W-2W^2
2W^2-60W+352=0
2(W^2-30W+176)=0
2(W-22)(W-8)=0
W-8=0
W=8 ANSWER.
THEN L=60-2*8
L=60-16
L=44 ANSWER.
PROOF
8*44=352
352=352